Exercise 181 (Part Level Submission) Gallagher Company applies manufacturing overhead to jobs on the basis of machine hours used. Overhead costs are expected to total $425,000 for the year, and machine usage is estimated at 125,000 hours. For the year, $450,000 of overhead costs are incurred and 130,000 hours are used. Compute the manufacturing overhead rate for the year. (Round answer to 2 decimal places, e.g. 52.75) Overhead rate per machine hour

Assuming the under- or overapplied overhead for the year is not allocated to inventory accounts, prepare the adjusting entry to assign the amount to cost of goods sold. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit
Answer #1

Overhead rate =Estimated overhead/Machine Hour

=$425000/125000=$3.4 /Hr rate

c

Overhead Applied =130000*3.4=$442000

Overhead Actual =$450000

Overhead Under applied =4420000-450000 =$8000

AGcuCSZDRP5XAAAAAElFTkSuQmCC